# DeepVoid: A Deep Learning Void Detector

### v1

@@ 3 Tidal Tensor Definition
Figure 2 shows the effect of changing the eigenvalue threshold in this scheme. All training and testing was done using tidal tensor based masks with $\lambda_{\mathrm{th}}=0.65$. We choose this value because it segments space into proportions of void, wall, filament, and halo morphologies that roughly match other methods. Given this threshold, TNG300 is composed of approximately 65% void, 26% wall, 9% filament, and 0.5% halo voxels.
4 Application of Deep Learning
@@
Our choice of CNNs for void-finding is driven by their success in segmenting volume-filling structures in other three-dimensional data, particularly in analysis of biomedical imaging (e.g., Niyas et al., 2022).
CNNs employ convolutional layers that apply trainable filters to the input data, enabling the identification of spatially localized features. Additionally, through successive layers and pooling operations, CNNs can capture complex and nonlinear features. The architecture we use in this work is a U-Net, a type of CNN based on the model of Ronneberger et al. (2015), originally designed to perform segmentation of neuronal structures imaged using electron microscopy. The U-Net architecture is named after the distinctive U-shaped network structure consisting of an encoding and decoding side. The encoder pathway, consisting of convolutional and pooling layers, progressively extracts higher-level features, enabling the network to learn a rich hierarchical representation of the input image. The decoder pathway, through upsampling and convolutional layers, recovers spatial details and combines them with corresponding features from the encoder pathway using concatenations. This facilitates the accurate localization of objects or regions in the segmentation output. For the problem of segmenting large-scale structure in the galaxy distribution, this neural network architecture allows us to train models that recognize the spatial patterns of voids, walls, filaments, and halos that gravity imprints on the galaxy distribution. Figure 3 illustrates the U-net architecture.
Figure 3: Schematic of a U-Net Architecture with three levels of resolution (‘depth’). In a), chunks of a density field are fed through the successive layers in b), with cubes representing 3D feature maps colored orange on the encoding side, and green on the decoding side. Numbers above the cubes indicate the number of filters in each layer. Colored arrows indicate operations performed between different layers. Horizontal gray arrows represent the merging of feature maps used to transfer small-scale spatial information from the encoding layers to the decoding layers. Without these concatenations, the U-net architecture is identical to that of an autoencoder. The predicted structural segmentation is shown in c).
@@
The activation function we employ is the Rectified Linear Unit (ReLU) function: $\phi(x)=x$ if $x\geq 0$ and 0 if $x<0$ (Glorot et al., 2011). ReLUs allow for extremely efficient computation, as there are no arithmetic operations necessary. Using ReLU after all convolutions (except the last) enables the network to cheaply incorporate non-linearity. MaxPooling layers in the CNN add non-linear behavior as well, since they only pass down the maximum value of $\phi(x)$ from a compact set of pixels. We use MaxPooling3D layers with a kernel size of (2,2,2), a stride of 2, and “same” padding to ensure that the output maintains the same shape as the input. However, since the stride is equal to the kernel size, the spatial dimensions are effectively reduced by a factor of two after each pooling operation. At the same time, the number of features channels doubles, allowing the network to learn progressively richer representations at increasingly coarser spatial scales.
@@
Overfitting is an inherent challenge in training deep neural networks. Given a finite dataset, almost any sufficiently complex model will eventually learn not only meaningful patterns but also noise and random fluctuations specific to the training data, reducing its ability to generalize to new inputs. This issue becomes apparent when the training loss continues to decrease while the validation loss plateaus or increases, as seen in Figure 4. To mitigate overfitting, we employ several strategies. First, we augment our dataset in the form of $90\degree$ rotations, where each subcube is rotated three times, effectively increasing the dataset size by a factor of four. Additionally, we split the dataset into training and validation sets, ensuring that we are evaluating model performance on data that it has not seen before. We also experiment with different learning rates to balance convergence speed and model stability. We found that a learning rate of 0.0003 allowed the models to converge efficiently while still enabling effective learning. We employ a learning rate scheduler that quarters the rate if the validation loss has not improved in 15 epochs, allowing the models to refine their parameters more gradually, improving generalization. Finally, we save the model weights corresponding to the lowest validation loss during training, which ensures that we retain the best performing model rather than one that may have overfit in later epochs.
4.1 U-Net Architecture
For our specific choice of CNN, we select the U-net architecture of Ronneberger et al. (2015). Similar to an autoencoder, it learns an internal compressed representation of the input as the information cascades down the encoding path. The data is then reconstructed by the network as the spatial resolution increases on its way up the decoding path. However, autoencoders suffer from a lack of locality. The U-net architecture addresses this by concatenating small-scale features from the encoding path to the decoding path. This allows the network to localize better while parsing data on multiple scales and has led to the U-net becoming one of the standard architectures for image processing networks. Throughout the rest of this paper, we will refer to the number of encoding/decoding layers as the ‘depth’ of the U-net, and the number of filters in the first convolutional block as the number of ‘filters’ for that particular DeepVoid model (see Figure 3 for reference).
@@ 4.1 U-Net Architecture
As information propagates through the U-net, undergoing convolutions, poolings, upsamplings, and concatenations, it finally converges at the final convolutional layer. This layer produces a feature tensor with $N_{\rm classes}$ as the last dimension. This tensor is then fed to the last activation function. Since we are performing a multi-label classification, we set the last activation layer to the Softmax function:
@@
where $\vec{z}$ is the input vector with components $(z_{0},\ldots z_{K})$ and $K$ is the number of classes (Bridle, 1990). The term in the denominator ensures that the probabilities of different classes for a pixel, or in our case a voxel, always add up to one. This ensures that the output of the Softmax function resembles a probability distribution over $K$ different possible outcomes. Strictly speaking, the Softmax outputs are not probabilities, but rather are more like normalized confidence scores (Guo et al., 2017). Finally, in order to predict a single class for each voxel, the Argmax function is used, which simply assigns the class with the greatest confidence score to each respective voxel.
4.2 Choosing a loss function
Once a model makes its predictions, a loss function is necessary to determine how well the predictions match the ground truth. In order to train a network, we need a differentiable loss function to minimize through backpropagation. The goal of training is to optimize the layer weights and biases such that it results in the lowest possible loss. The ‘loss landscape’ is the high dimensional space of those parameters, which is typically conceptualized as a surface. Ideally, the training results in the model finding a global minimum, i.e., as well as the model can perform given its hyperparameters. However, in practice, the optimizer is really trying to find the best possible local minimum in that landscape.
@@ 4.2 Choosing a loss function
$L_{\rm CCE}(y,\hat{y})=-\sum_{i}^{N}y_{i}\log(\hat{y_{i}}),$ (6)
where $y$ and $\hat{y}$ are vectors denoting the actual and predicted labels respectively. The ‘sparse’ part of SCCE indicates that the labels are encoded as integers instead of one-hot vectors, which greatly reduces the memory overhead during training.
@@ 5 Training and Validation
Training in this work was performed on two high-performance computing clusters: Picotte at the Drexel University Research Computing Facility and Nautilus at the National Research Platform. Each training run on Picotte used two NVIDIA Tesla V100 GPUs, each with 32 GB of memory, while predictions were performed on one Tesla V100 GPU. Trainings on Nautilus used various configurations of GPUs. We construct our 3D U-nets using the Keras library (Chollet et al., 2015) that runs on top of the TensorFlow library (Abadi et al., 2015) in Python. Additionally, we use the Adam (adaptive moment estimation, Kingma & Ba, 2014) stochastic gradient descent algorithm with an initial learning rate of 0.0003 and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-7}$.
5.1 Construction of training data
@@
We construct training data by first splitting the density and mask fields into smaller, overlapping subcubes. Each subcube overlaps the previous one by half. TNG’s $512^{3}$ grid is broken into 343 $128^{3}$ overlapping subcubes. For data augmentation purposes and to help avoid overfitting, each subcube is rotated 90° three times, to account for the lack of rotational invariance in CNNs. Therefore, the TNG dataset consists of 1372 subcubes. The resulting subcubes are then split into training and validation sets in an 80%/20% split. The subcubes are also shuffled during training to avoid feeding the network batches of consecutive rotated subcubes. Note that all metrics reported in this work are computed using the validation set.
@@
Prior to training, each mass density field is scaled (‘min-maxed’) to the range $[0,1]$. This scaling allows our models’ training to converge quicker and without any extreme gradients that could lead to the model failing to learn. If the input features are not scaled, some gradients calculated in the course of backpropagation can ‘explode’, becoming excessively large and hindering further training. Similarly, gradients can also ‘vanish’, which results in slow training and can often cause the model to get stuck in a suboptimal local minimum of the loss landscape. Deep neural networks such as the U-nets we employ are especially susceptible to these effects, because gradients are multiplied across many layers during the course of backpropagation (Goodfellow et al., 2016). We also examined using the logarithm of mass density, and standardization, a normalization technique common in computer vision, where the training data is normalized such that its mean is zero and standard deviation is one. We found minimal differences in predictive performance between min-max normalization and standardization and elected to use min-max.
@@
Figure 4: Metrics during the training of a DeepVoid base model. Top: SCCE loss vs. training epochs. Middle: Matthews correlation coefficient for training and testing datasets during training. Note that overfitting begins to occur around epoch 40, where the training and validation scores begin to diverge. Bottom: learning rate with reductions by a quarter every time the validation loss has not improved for 10 epochs.
5.2 Training
@@
Each DeepVoid network was set to train for a maximum of 300 epochs, halting if the validation loss of the model has not improved in 25 epochs, and we save the weights that resulted in the lowest validation loss, as seen in Figure 4. Early stopping is one of the chief ways that we can avoid overfitting when training models on a relatively small dataset. The average number of epochs is 85 across all trainings, with the maximum being 250. To avoid overshooting potentially better local minima in the loss landscape, we also set the learning rate to quarter if the validation loss has not improved in 15 epochs. This adjustment of learning rate can also help the model fine-tune its weights more delicately in the final stage of training when the validation loss has leveled off, leading to better predictive performance and generalization.
5.3 Curriculum learning
@@
Curriculum learning (CL) is a training strategy used in machine learning that involves beginning with ‘easier’ tasks and progressively increasing the tasks’ difficulty (Bengio et al., 2009; Soviany et al., 2021). The idea is to have the model establish a strong foundation before attempting to classify more difficult examples. In the method of transfer learning, training of the model for the harder task begins in a location in the loss landscape that it inherits from the model trained on the easier task, then further training fine-tunes the model for the harder task. The goal of DeepVoid is to train a model that classifies regions of space based on the spatial distribution of galaxies, which act as tracers of the matter distribution. In this case a model might first be trained on a relatively dense data set (the easy task), then the weights from that model are used as the initial conditions for training a model that performs well on a sparse data set (the hard task).
@@
We initially train ‘base’ models on a mass density field computed by considering all DM particles in the simulation. For TNG300-3-Dark, this results in an average interparticle separation of $\lambda=0.33$ $h^{-1}$ Mpc. To obtain sparser samples that more closely resemble the actual distribution of galaxies that we observe at medium to high redshifts, we subsample each simulation’s halos/subhalos. To do so, we take the most massive $N_{\rm subhalos}$ such that the intertracer separation is the desired value. Next, we apply the CIC interpolation scheme to the subhalos to obtain the subhalo matter density contrast as described in Section 3. The multiclass truth table remains the same, as it was computed from the gravitational potential of the full density field. This allows the model to train and develop weights that can take a sparse density field as input and predict structural classifications that match the labels computed from the full density field. We created density volumes with average intertracer separations of $\lambda=[1,3,5,7,10]$ $h^{-1}$ Mpc, with most of our tests evaluating performance on $\lambda=10$ $h^{-1}$ Mpc.
5.4 Prediction

### v2

@@ III Tidal Tensor Definition
Figure 2 shows the effect of changing the eigenvalue threshold in this scheme. All training and testing wer done using tidal-tensor-based masks with $\lambda_{\mathrm{th}}=0.65$. We choose this value because it segments space into proportions of void, wall, filament, and halo morphologies that roughly match other methods. Given this threshold, TNG300 is composed of approximately 65% void, 26% wall, 9% filament, and 0.5% halo voxels.
IV Application of Deep Learning
IV.1 Why Use Deep Learning?
Here, we describe the deep learning method that we adopt for translating the tidal tensor classification of large-scale structure into a method that can be applied to sparse galaxy survey data. As described above in Section III, if we were given the exact mass distribution in the Universe, as is true for cosmological simulations, then there would be no need for deep learning; one would simply compute the potential and its derivatives to form the tidal tensor. In the first application described below (Section VI.1), we build a deep learning model that is trained on the full DM density field and demonstrate that it accurately recovers the structural classes of the truth table that was also computed from that full DM distribution. In that specific test case, DeepVoid could be described as an “emulator” of the structural class definitions that were computed from the tidal tensor. The need for deep learning arises when we move to training models appropriate for analysis of sparse galaxy data (Section VII). In this significantly more challenging application, the model must learn to recognize voids, filaments, walls, and halos from their spatial patterns imprinted on the distribution of these sparse and biased tracers.
IV.2 Convolutional Neural Networks
@@
Our choice of CNNs for void-finding is driven by their success in segmenting volume-filling structures in other three-dimensional data, particularly in the analysis of biomedical imaging (e.g., Niyas et al., 2022).
CNNs employ convolutional layers that apply trainable filters to the input data, enabling the identification of spatially localized features. Additionally, through successive layers and pooling operations, CNNs can capture complex and nonlinear features.
@@
The activation function we employ is the rectified linear unit (ReLU) function: $\phi(x)=x$ if $x\geq 0$ and 0 if $x<0$ (Glorot et al., 2011). ReLUs allow for extremely efficient computation, as there are no arithmetic operations necessary. Using ReLU after all convolutions (except the last) enables the network to cheaply incorporate nonlinearity. MaxPooling layers in the CNN add nonlinear behavior as well, since they only pass down the maximum value of $\phi(x)$ from a compact set of pixels. We use MaxPooling3D layers with a kernel size of (2,2,2), a stride of 2, and “same” padding to ensure that the output maintains the same shape as the input. However, since the stride is equal to the kernel size, the spatial dimensions are effectively reduced by a factor of 2 after each pooling operation. At the same time, the number of feature channels doubles, allowing the network to learn progressively richer representations at increasingly coarser spatial scales.
@@
Overfitting is an inherent challenge in training deep neural networks. Given a finite dataset, almost any sufficiently complex model will eventually learn not only meaningful patterns but also noise and random fluctuations specific to the training data, reducing its ability to generalize to new inputs. This issue becomes apparent when the training loss continues to decrease while the validation loss plateaus or increases, as seen in Figure 4. To mitigate overfitting, we employ several strategies. First, we augment our dataset in the form of $90\degree$ rotations, where each subcube is rotated three times, effectively increasing the dataset size by a factor of 4. Additionally, we split the dataset into training and validation sets, ensuring that we are evaluating model performance on data that it has not seen before. We also experiment with different learning rates to balance convergence speed and model stability. We found that a learning rate of 0.0003 allowed the models to converge efficiently while still enabling effective learning. We employ a learning rate scheduler that quarters the rate if the validation loss has not improved in 15 epochs, allowing the models to refine their parameters more gradually, improving generalization. Finally, we save the model weights corresponding to the lowest validation loss during training, which ensures that we retain the best-performing model rather than one that may have overfitted in later epochs.
Figure 3: Schematic of a U-Net architecture with three levels of resolution (‘depth’). In (a), chunks of a density field are fed through the successive layers in (b), with cubes representing 3D feature maps colored orange on the encoding side, and green on the decoding side. Numbers next to the curly brackets indicate the size of the activation images in a given layer, while numbers next to the vertical ellipses indicate the number of filters in each layer. Colored arrows indicate operations performed between different layers. Horizontal gray arrows represent the merging of feature maps used to transfer small-scale spatial information from the encoding layers to the decoding layers. Without these concatenations, the U-Net architecture is identical to that of an autoencoder. The predicted structural segmentation is shown in (c).
IV.3 U-Net Architecture
The architecture we use in this work is a U-Net, a type of CNN based on the model of Ronneberger et al. (2015), originally designed to perform segmentation of neuronal structures imaged using electron microscopy. Similar to an autoencoder, the U-Net learns an internal compressed representation of the input as the information cascades down the encoding path. The data is then reconstructed by the network as the spatial resolution increases on its way up the decoding path. However, autoencoders suffer from a lack of locality. The U-net architecture addresses this by concatenating small-scale features from the encoding path to the decoding path. This allows the network to localize better while parsing data on multiple scales and has led to the U-net becoming one of the standard architectures for image processing networks. The U-Net architecture is named after the distinctive U-shaped network structure consisting of an encoding and decoding side. The encoder pathway, consisting of convolutional and pooling layers, progressively extracts higher-level features, enabling the network to learn a rich hierarchical representation of the input image. The decoder pathway, through upsampling and convolutional layers, recovers spatial details and combines them with corresponding features from the encoder pathway using concatenations. This facilitates the accurate localization of objects or regions in the segmentation output. For the problem of segmenting large-scale structure in the galaxy distribution, this neural network architecture allows us to train models that recognize the spatial patterns of voids, walls, filaments, and halos that gravity imprints on the galaxy distribution. Figure 3 illustrates the U-net architecture (see He et al. (2019); Berger and Stein (2019); Jamieson et al. (2023) for other applications of CNNs and UNet to 3D datasets in astronomy).
Throughout the rest of this paper, we will refer to the number of encoding/decoding layers as the ‘depth’ of the U-net, and the number of filters in the first convolutional block as the number of ‘filters’ for that particular DeepVoid model (see Figure 3 for reference).
@@ IV.3 U-Net Architecture
As information propagates through the U-net, undergoing convolutions, poolings, upsamplings, and concatenations, it finally converges at the final convolutional layer. This layer produces a feature tensor with $N_{\rm classes}$ as the last dimension. This tensor is then fed to the last activation function. Since we are performing a multilabel classification, we set the last activation layer to the Softmax function:
@@
where $\vec{z}$ is the input vector with components $(z_{0},\ldots z_{K})$ and $K$ is the number of classes (Bridle, 1990). The term in the denominator ensures that the probabilities of different classes for a pixel, or in our case, a voxel, always add up to one. This ensures that the output of the Softmax function resembles a probability distribution over $K$ different possible outcomes. Strictly speaking, the Softmax outputs are not probabilities, but rather are more like normalized confidence scores (Guo et al., 2017). Finally, in order to predict a single class for each voxel, the Argmax function is used, which simply assigns the class with the greatest confidence score to each respective voxel.
IV.4 Choosing a loss function
Once a model makes its predictions, a loss function is necessary to determine how well the predictions match the ground truth. In order to train a network, we need a differentiable loss function to minimize through backpropagation. The goal of training is to optimize the layer weights and biases such that it results in the lowest possible loss. The ‘loss landscape’ is the high-dimensional space of those parameters, which is typically conceptualized as a surface. Ideally, the training results in the model finding a global minimum, i.e., as well as the model can perform given its hyperparameters. However, in practice, the optimizer is really trying to find the best possible local minimum in that landscape.
@@ IV.4 Choosing a loss function
[…]
where $y$ and $\hat{y}$ are vectors denoting the actual and predicted labels, respectively. The ‘sparse’ part of SCCE indicates that the labels are encoded as integers instead of one-hot vectors, which greatly reduces the memory overhead during training.
@@ V Training and Validation
Training in this work was performed on two high-performance computing clusters: Picotte at the Drexel University Research Computing Facility and Nautilus at the National Research Platform. Each training run on Picotte used two NVIDIA Tesla V100 GPUs, each with 32 GB of memory, while predictions were performed on one Tesla V100 GPU. Training on Nautilus used various configurations of GPUs. We construct our 3D U-nets using the Keras library (Chollet and others, 2015) that runs on top of the TensorFlow library (Abadi et al., 2015) in Python. Additionally, we use the Adam (adaptive moment estimation, Kingma and Ba, 2014) stochastic gradient descent algorithm with an initial learning rate of 0.0003 and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-7}$.
V.1 Construction of training data
@@
We construct training data by first splitting the density and mask fields into smaller subcubes. TNG’s $512^{3}$ grid is broken into 343 $128^{3}$ subcubes. For data augmentation purposes and to help avoid overfitting, each subcube is rotated 3 times by 90°  to take advantage of the lack of rotational invariance in CNNs. Therefore, the TNG dataset consists of 1372 subcubes. The resulting subcubes are then split into training and validation sets in an 80%/20% split. The subcubes are also shuffled during training to avoid feeding the network batches of consecutive rotated subcubes.
Note that all metrics reported in this work are computed using the validation dataset. We use the same validation data set to compute performance metrics when choosing hyperparameters, when computing validation loss during training, and for computing final performance metrics. The fact that we do not use a separate testing dataset to evaluate final performance metrics is of possible concern if the examples in the validation dataset are not representative of the space of possible structures. That would also be true for a small separate testing dataset. Thus, we use a single validation/testing dataset to include a large set of different examples from the TNG300 simulation box. If we are unlucky and the examples in the validation data are quite dissimilar from the training data (assuming here that the training dataset is large enough to be representative), then this could bias both the “stopping” point of training and the performance metrics: validation loss would exceed training loss too soon, causing an early stop to training, and result in poor performance metrics computed on the validation data, both because the training was stopped too soon and because the validation data are dissimilar to the training data. Splitting TNG300 into smaller, separate training/validation/testing datasets would increase the probability of such a result. Another possible concern is that using the same validation data to set hyperparameters admits the possibility that our choices of hyperparameters affect the stopping point of training (based on validation loss) and vice versa.
We note that while TNG300 may be considered a medium-sized box for cosmological studies, it still should contain a few thousand voids (assuming a mean void size of $\sim 10$Mpc; Platen et al. (2008)) and the subcube rotation done during data augmentation effectively increases the training volume by a factor of 4. One could add more consecutive rotations and also mirror operations to further increase the size of the training dataset without risking overfitting, given the lack of rotational invariance of CNNs.
@@
Prior to training, each mass density field is scaled (“min-maxed”) to the range $[0,1]$. The min-max operation is applied to the final snapshot. In the case of processing different snapshots one could either compute a global min-max for all the snapshots used during training, or one could apply the min-max range computed from one snapshot to the rest. This scaling allows our models’ training to converge more quickly and without any extreme gradients that could lead to the model failing to learn. If the input features are not scaled, some gradients calculated in the course of backpropagation can “explode,” becoming excessively large and hindering further training. Similarly, gradients can also “vanish,” which results in slow training and can often cause the model to get stuck in a suboptimal local minimum of the loss landscape. Deep neural networks such as the U-Nets we employ are especially susceptible to these effects, because gradients are multiplied across many layers during the course of backpropagation (Goodfellow et al., 2016). We also examined using the logarithm of mass density, and standardization, a normalization technique common in computer vision, where the training data is normalized such that its mean is zero and standard deviation is one. We found minimal differences in predictive performance between min-max normalization and standardization and elected to use min-max.
@@
Figure 4: Metrics during the training of a DeepVoid base model. Upper: SCCE loss vs. training epochs. Lower: Matthews correlation coefficient for training and testing datasets during training. Note that while around epoch 40 the training and validation scores begin to diverge we further monitor the validation loss to stop training when it does not improve for 25 epochs.
V.2 Training
@@
Each DeepVoid network was set to train for a maximum of 300 epochs, halting if the validation loss of the model had not improved in 25 epochs, and we saved the weights that resulted in the lowest validation loss, as seen in Figure 4. Early stopping is one of the chief ways that we can avoid overfitting when training models on a relatively small dataset. The average number of epochs is 85 across all trainings, with the maximum being 250. To avoid overshooting potentially better local minima in the loss landscape, we set the learning rate to quarter if the validation loss has not improved in 15 epochs. This adjustment of learning rate can also help the model fine-tune its weights more delicately in the final stage of training when the validation loss has leveled off, leading to better predictive performance and generalization.
V.3 Curriculum learning
@@
Curriculum learning (CL) is a training strategy used in machine learning that involves beginning with ‘easier’ tasks and progressively increasing the tasks’ difficulty (Bengio et al., 2009; Soviany et al., 2021). The idea is to have the model establish a strong foundation before attempting to classify more difficult examples. In the method of transfer learning, training of the model for the harder task begins in a location in the loss landscape that it inherits from the model trained on the easier task, then further training fine-tunes the model for the harder task. The goal of DeepVoid is to train a model that classifies regions of space based on the spatial distribution of galaxies, which act as tracers of the matter distribution. In this case, a model might first be trained on a relatively dense dataset (the easy task), then the weights from that model are used as the initial conditions for training a model that performs well on a sparse data set (the hard task).
@@
We initially train ‘base’ models on a mass density field computed by considering all DM particles in the simulation. For TNG300-3-Dark, this results in an average interparticle separation of $\lambda=0.33$ $h^{-1}$ Mpc. To obtain sparser samples that more closely resemble the actual distribution of galaxies that we observe at medium to high redshifts, we subsample each simulation’s halos/subhalos. To do so, we take the most massive $N_{\rm subhalos}$ such that the intertracer separation is the desired value. Next, we apply the CIC interpolation scheme to the subhalos to obtain the subhalo matter density contrast as described in Section III. The multiclass truth table remains the same, as it was computed from the gravitational potential of the full density field. This allows the model to train and develop weights that can take a sparse density field as input and predict structural classifications that match the labels computed from the full density field. We created density volumes with average intertracer separations of $\lambda=[1,3,5,7,10]$ $h^{-1}$ Mpc, with most of our tests evaluating performance on $\lambda=10$ $h^{-1}$ Mpc.
V.4 Prediction